# An Analysis of Concept Bottleneck Models: Measuring, Understanding, and Mitigating the Impact of Noisy Annotations

### v1

@@ 5 Conclusion
Further analyses Beyond the primary experiments, we undertook a suite of supplementary studies to deepen our evaluation of noise robustness in CBMs. First, we performed ablation studies comparing alternative training strategies across varying noise levels. Second, we benchmarked our framework against representative noise-mitigation baselines. Third, we evaluated robustness on several advanced CBM variants under diverse noise configurations. Finally, we carried out auxiliary experiments probing orthogonal aspects of resilience. Comprehensive results are available in Appendix D.
@@
Comparisons Prior work has pursued robustness in related settings, yet each tackles a distinct aspect. Sinha et al. [43] analyzed adversarial perturbations of continuous concepts to develop defenses against malicious attacks, Sheth and Ebrahimi Kahou [41] enforced disentangled representations via auxiliary losses to improve generalization under distribution shifts, and Penaloza et al. [29] proposed a preference-optimization objective to reduce noise sensitivity. In contrast, our study offers the first unified framework that measures, explains, and mitigates noise effects through the combined use of sharpness-aware training and uncertainty-guided correction. Theoretical analysis and extensive empirical evaluation substantiate the effectiveness of this unified approach.
@@
Limitations  Despite the substantial robustness gains provided by our framework, several limitations warrant further investigation. First, the method assumes well-calibrated uncertainty estimates, which is an assumption that may not hold in data-scarce or highly imbalanced regimes. Second, the reliance on a linear target predictor, while beneficial for interpretability, may hinder generalization on tasks with more complex decision boundaries. Third, our study is confined to binary concept labels; extending the framework to hierarchical, multi-class, or continuous concepts remains a promising avenue for future work. Addressing these limitations would further enhance the practical utility of CBMs in real-world environments characterized by noisy or incomplete supervision.
@@
Future work Our research opens up new avenues for future research. First, extending our mitigation strategy to hierarchical, multi-class, or continuous concept spaces [31, 28] would enable applications such as fine-grained scene understanding and clinical severity grading. Second, the wider adoption of concept labels generated by large language models [58, 26, 44, 45] calls for a systematic study of their noise properties and their downstream impact on CBMs, which may offer important insights for both annotation protocols and model design. Third, incorporating semi-supervised objectives [15] and using active selection to identify the most informative concepts can decrease reliance on expensive or error-prone human labeling. Progress in these areas will further advance the deployment of robust and interpretable CBMs in real-world scenarios where supervision is noisy or incomplete.
@@
    Mitigating effects We propose a two-stage strategy: (i) sharpness-aware minimization during training, which stabilizes noise-sensitive concepts; and (ii) uncertainty-guided intervention at inference, which corrects only the most unreliable concept predictions. Both components are theoretically motivated and supported by extensive empirical evidence.
Collectively, these contributions establish a principled framework for building concept-based models that remain reliable and transparent in the presence of noisy supervision.
@@ Appendix A Theoretical Analysis for Sharpness-Aware Minimization for CBMs
[…]
@@ Appendix B Theoretical Justification for Uncertainty-Based Intervention
Let $\widehat{c}=(\widehat{c}_{1},\dots,\widehat{c}_{k})$ denote the vector of predicted concept values produced by a CBM, and let $c^{*}=(c_{1}^{*},\dots,c_{k}^{*})$ denote the corresponding ground-truth concept labels. Let $f$ be the target predictor and $\ell$ the task loss function.
@@
$\ell(f(\widehat{c}_{-i}),y)$ $\approx\ell(f(\widehat{c}),y)+\nabla_{\widehat{c}}\ell(f(\widehat{c}),y)^{\top}(\widehat{c}_{-i}-\widehat{c})$ (18)
@@
Moreover, suppose that the predicted value $f(\widehat{c})$ lies within a bounded interval $[a,b]\subset\mathbb{R}$. Then, for any reference point $f_{0}\in[a,b]$, it follows from Lipschitz continuity that:
@@
To isolate the effect of uncertainty, we assume $|w_{i}|$ is approximately constant across all concepts. Under this simplification and B.5:
@@
Under B.4 and B.5, and assuming the linear-noise model (28), the uncertainty-based top-$s$ set $\mathcal{U}$ recovers the susceptibility-based set $\mathcal{S}$ in the low-noise regime:
@@
By Lemma B.8, the rankings $\pi_{u}$ and $\pi_{\delta}$ converge in probability as the noise variance $\sigma^{2}$ tends to zero. Since both $\mathcal{U}$ and $\mathcal{S}$ are determined by the top-$s$ indices in their respective rankings, their equality follows in the limit. ∎
This analysis provides both intuitive and formal support for the use of uncertainty-guided interventions in CBMs. While the assumptions (e.g., constant weight or the approximate linear relationship between entropy and prediction error) may not hold strictly in practice, empirical evidence (see Figure 10) demonstrates a consistent and positive correlation between predictive uncertainty and susceptibility, validating the practical utility of this approach.
@@ C.2 CBM Training Strategies
Here, $\ell_{C_{j}}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ denotes the loss function for the $j$-th concept, and $\ell_{Y}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ denotes the loss for target prediction. Our main experiments are conducted using the Independent strategy, and comparative results for the other strategies are provided in Section D.1.
@@ C.3 Implementation Details
For concept prediction, we fine-tune InceptionV3 [46], pre-trained on ImageNet [6], on the CUB [52] and AwA2 [56] datasets. The target predictor $f$ is implemented as a single-layer linear classifier, following the standard setup in Koh et al. [18]. All experiments are conducted with three random seeds, and we report the average performance. Training and evaluation are conducted on NVIDIA GeForce RTX 3090 (24GB), RTX A6000 (48GB), and A100 (80GB) GPUs.
@@
All additional CBM variants shown in Section D.2 are trained on the CUB dataset for 300 epochs using the Adam optimizer [17] with a fixed learning rate of $10^{-4}$, following Vandenhirtz et al. [50]. For the independently trained autoregressive CBM [13], training is scheduled such that two-thirds of the total epochs are allocated to learning the concept predictor $g$, and the remaining one-third to the target predictor $f$, consistent with Vandenhirtz et al. [50]. We employ ResNet-18 [14] as the backbone for CEM, SCBM, and autoregressive CBM, and ResNet-101 for ECBM. The batch size is set to 256 for CEM, SCBM, and autoregressive CBM models, and 64 for ECBM. When applying SAM to these models, we perform a grid search over $\rho\in\{0.01,0.05,0.1\}$, while keeping all other training hyper-parameters aligned with those used for the Adam baseline.
We conduct a targeted hyperparameter search for each label noise mitigation technique evaluated in Section D.3, including label smoothing regularization (LSR) [46], symmetric cross-entropy learning (SL) [54], and SAM. For each method, we tune the corresponding parameters (see Table 4), such as the smoothing factor $\epsilon$ for LSR, regularization coefficients $\alpha$ and $\beta$ for SL, and the sharpness radius $\rho$ for SAM, under varying levels of noise ($\gamma\in\{0.0,0.1,0.2,0.3,0.4\}$). The best hyperparameters are selected based on task accuracy.
@@ C.4 Label Noise Injection
    Symmetric noise. Each class label is randomly flipped to one of the remaining $N-1$ classes with uniform probability. Specifically, for a dataset with $N$ classes, the probability of mislabeling a sample as any incorrect class is $\frac{1}{N-1}$. This simulates class-independent, uniformly distributed noise. An analogous strategy is applied to concepts, where each concept label is flipped uniformly to a different one.
@@
Symmetric noise captures general annotation uncertainty, asymmetric noise models systematic labeling biases, and grouped noise reflects domain-driven confusions within related concept groups. Together, these diverse noise settings constitute a comprehensive testbed for assessing the robustness of concept-based models. Experimental results under these conditions are presented in Section D.4.
@@ C.5 Concept Selection Criteria
This section provides a detailed explanation of the concept selection criteria, extensively analyzed in previous work [42]. We present three criteria (Random, UCP, CCTP) addressed in the main paper, followed by additional criteria (LCP, ECTP, EUDTP) discussed in Section D.8.
@@
    Expected change in target prediction (ECTP). This criterion prioritizes concepts whose correction results in the greatest expected change in the target predictive distribution. Scores are defined as $s_{i}=(1-\widehat{c}_{i})D_{\mathrm{KL}}(\widehat{y}_{\widehat{c}_{i}=0}\|\widehat{y})+\widehat{c}_{i}D_{\mathrm{KL}}(\widehat{y}_{\widehat{c}_{i}=1}\|\widehat{y})$ where $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence, and $\widehat{y}_{\widehat{c}_{i}=0},\widehat{y}_{\widehat{c}_{i}=1}$ denote target predictions after interventions. The intuition is to intervene on concepts whose rectification significantly alters the target prediction [39].
@@
    Expected uncertainty decrease in target prediction (EUDTP). Concepts that lead to the largest expected reduction in the entropy of the target predictive distribution upon intervention are selected. Formally, scores are defined as $s_{i}=(1-\widehat{c}_{i})\mathcal{H}(\widehat{y}_{\widehat{c}_{i}=0})+\widehat{c}_{i}\mathcal{H}(\widehat{y}_{\widehat{c}_{i}=1})-\mathcal{H}(\widehat{y})$. This method favors concepts that substantially reduce target prediction uncertainty when corrected [11].
@@ D.1 Effect of Training Strategies under Noise
CBMs can be optimized with three training paradigms: independent, sequential, and joint (see Section C.2 for details). We evaluate their robustness by reporting task-level and concept-level accuracy across different noise ratios $\gamma$. Figure 12 presents task accuracy across training strategies, with accompanying radar plots visualizing noise tolerance; larger and more regular polygons reflect greater robustness to noise.
Every training paradigm suffers noticeable degradation as label corruption intensifies. On AwA2, for instance, the joint model starts with the best clean accuracy but still drops from 88.8% to 81.7% when 40% of the labels are noisy. The independent and sequential variants fare even worse and largely collapse on CUB under the same noise level. Table 5 further reveals declines at the concept level. Although the joint paradigm better preserves task-level performance, its concept accuracy plunges below 75% once the noise ratio rises above roughly 30% on CUB and exceeds about 10% on AwA2. Because the joint objective simultaneously minimizes task and concept losses, it compromises concept fidelity as label quality declines. These findings corroborate earlier evidence of a performance–interpretability tradeoff [35] and underscore the need for training methods that preserve the robustness of CBM under noisy supervision.
@@ D.2 Robustness of CBM Variants under Noisy Supervision
We also investigate the effects of noise on several CBM variants, including CEM [59], ECBM [57], autoregressive CBM [13], and SCBM [50], trained on the CUB dataset with noise levels $\gamma\in[0.0,0.4]$. Figure 13 presents task accuracy, while Table 6 reports concept prediction accuracy under both the standard Adam optimizer (Base) and sharpness-aware minimization (SAM).
@@
D.3 Comparative Study of Noise Mitigation Methods
@@
Concept Acc 95.9$\pm$$0.1$ 91.3$\pm$$0.0$ 86.5$\pm$$0.2$ 70.7$\pm$$7.7$
LSR Task Acc 72.8$\pm$$0.4$ 60.4$\pm$$0.4$ 46.8$\pm$$1.6$ 15.0$\pm$$12.9$
Concept Acc 95.0$\pm$$0.1$ 91.5$\pm$$0.1$ 86.6$\pm$$0.2$ 81.6$\pm$$0.2$
SL Task Acc 70.6$\pm$$1.3$ 60.5$\pm$$0.4$ 48.0$\pm$$1.7$ 27.0$\pm$$7.5$
Concept Acc 97.3$\pm$$0.0$ 95.1$\pm$$0.1$ 92.7$\pm$$0.1$ 90.0$\pm$$0.1$
SAM Task Acc 79.7$\pm$$0.1$ 67.7$\pm$$0.6$ 50.4$\pm$$2.0$ 26.6$\pm$$1.2$
Table 7 summarizes task and concept accuracy across noise levels. Across all conditions, SAM consistently outperforms LSR and SL. Under clean supervision, SAM achieves a task accuracy of 79.7%, surpassing LSR (72.8%) and SL (70.6%). Its advantages are more pronounced under noise: at 20% and 30% corruption, SAM outperforms the second-best method by 6.1% and 8.4% in concept accuracy, respectively. These results demonstrate the robustness of SAM in noisy settings. While our evaluation focuses on selected methods, other techniques (e.g., co-teaching [47] or MentorNet [16]) offer promising alternatives through sample selection and teacher-student training. Integrating such techniques with CBMs represents a valuable direction for future work. Overall, these results position SAM as an effective and scalable approach for robust concept-based learning, as further supported by our theoretical insights in Appendix A.
(a) Task accuracy
(b) Concept accuracy
D.4 Robustness under Diverse Noise Scenarios
We further investigate CBM performance under realistic noise scenarios, specifically asymmetric [12] and grouped noise, using the CUB dataset. Asymmetric noise cyclically mislabels categories (e.g., confusing similar animal species), while grouped noise randomly corrupts labels within similar concept groups (e.g., wing color) as discussed in Section C.4.
As shown in Figure 14, CBMs experience substantial performance degradation under both noise types, broadly consistent with trends observed under symmetric noise. While both asymmetric and grouped noise lead to slightly more pronounced declines in concept accuracy at higher corruption levels (e.g., 40%), the overall performance patterns remain comparable. These findings indicate that CBMs are generally vulnerable to various forms of label noise, reinforcing the need for robustness strategies that extend beyond simple symmetric corruption and account for more realistic noise scenarios.
D.5 Evaluating Concept Noise in LLM-Generated Annotations

### v2

@@ 6 Conclusion
Further analyses Beyond the primary experiments, we undertook a suite of supplementary studies to deepen our evaluation of noise robustness in CBMs. First, we performed ablation studies comparing alternative training strategies across varying noise levels. Second, we benchmarked our framework against representative noise‑mitigation baselines. Third, we evaluated robustness on several advanced CBM variants under diverse noise configurations. Finally, we carried out auxiliary experiments probing orthogonal aspects of resilience. Comprehensive results are available in Appendix˜D.
@@
Comparisons Prior work has pursued robustness in related settings, yet each tackles a distinct aspect. Sinha et al. [44] analyzed adversarial perturbations of continuous concepts to develop defenses against malicious attacks, Sheth and Ebrahimi Kahou [42] enforced disentangled representations via auxiliary losses to improve generalization under distribution shifts, and Penaloza et al. [30] proposed a preference-optimization objective to reduce noise sensitivity. In contrast, our study offers the first unified framework that measures, explains, and mitigates noise effects through the combined use of sharpness‑aware training and uncertainty‑guided correction. Theoretical analysis and extensive empirical evaluation substantiate the effectiveness of this unified approach.
@@
Limitations  Despite the substantial robustness gains provided by our framework, several limitations warrant further investigation. First, the method assumes well‑calibrated uncertainty estimates, which is an assumption that may not hold in data‑scarce or highly imbalanced regimes. Second, the reliance on a linear target predictor, while beneficial for interpretability, may hinder generalization on tasks with more complex decision boundaries. Third, our study is confined to binary concept labels; extending the framework to hierarchical, multi‑class, or continuous concepts remains a promising avenue for future work. Addressing these limitations would further enhance the practical utility of CBMs in real‑world environments characterized by noisy or incomplete supervision.
@@
Future work Our research opens up new avenues for future research. First, extending our mitigation strategy to hierarchical, multi-class, or continuous concept spaces [32, 29] would enable applications such as fine-grained scene understanding and clinical severity grading. Second, the wider adoption of concept labels generated by large language models [60, 27, 45, 46] calls for a systematic study of their noise properties and their downstream impact on CBMs, which may offer important insights for both annotation protocols and model design. Third, incorporating semi-supervised objectives [16] and using active selection to identify the most informative concepts can decrease reliance on expensive or error-prone human labeling. Progress in these areas will further advance the deployment of robust and interpretable CBMs in real‑world scenarios where supervision is noisy or incomplete.
@@
    Mitigating effects We propose a two‑stage strategy: (i) sharpness‑aware minimization during training, which stabilizes noise‑sensitive concepts; and (ii) uncertainty‑guided intervention at inference, which corrects only the most unreliable concept predictions. Both components are theoretically motivated and supported by extensive empirical evidence.
Collectively, these contributions establish a principled framework for building concept‑based models that remain reliable and transparent in the presence of noisy supervision.
@@ Appendix A Theoretical Analysis for Sharpness-Aware Minimization for CBMs
$\text{1-{SAM}: }\nabla_{w+\varepsilon}\ell(w+\varepsilon,x,t)=t\sigma(-tf(w+\varepsilon,x))\nabla_{w+\varepsilon}f(w+\varepsilon_{i},x)$
$\text{Jacobian {SAM}: }\Delta^{\mathrm{J\text{-}SAM}}\ell(w+\varepsilon,x,t)=t\sigma(-tf(w,x))\nabla_{w+\varepsilon}f(w+\varepsilon,x).$
@@ Appendix B Theoretical Justification for Uncertainty-Based Intervention
Let $\widehat{c}=(\widehat{c}_{1},\dots,\widehat{c}_{k})$ denote the vector of predicted concept values produced by a CBM trained on potentially noisy concept labels, and let $c^{*}=(c_{1}^{*},\dots,c_{k}^{*})$ denote the corresponding ground-truth concept labels. Let $f$ be the target predictor and $\ell$ the task loss function.
@@
Since $\widehat{c}$ encodes the consequences of training-time noise, $\delta_{i}$ quantifies how the model’s noise-induced uncertainty at inference translates into potential performance gains through targeted correction.
@@
[…]
@@
Moreover, suppose that the predicted value f(c^)f($\widehat${c}) lies within a bounded interval $a,b]\subset\mathbb{R}$. Then, for any reference point $f_{0}\in[a,b]$, it follows from Lipschitz continuity that:
@@
To isolate the effect of uncertainty, we assume $|w_{i}|$ is approximately constant across all concepts. Under this simplification and ˜B.5:
@@
Under ˜B.4 and B.5, and assuming the linear-noise model (28), the uncertainty-based top-$s$ set $\mathcal{U}$ recovers the susceptibility-based set $\mathcal{S}$ in the low-noise regime:
@@
By Lemma˜B.8, the rankings $\pi_{u}$ and $\pi_{\delta}$ converge in probability as the noise variance $\sigma^{2}$ tends to zero. Since both $\mathcal{U}$ and $\mathcal{S}$ are determined by the top-$s$ indices in their respective rankings, their equality follows in the limit. ∎
This analysis provides both intuitive and formal support for the use of uncertainty-guided interventions in CBMs. While the assumptions (e.g., constant weight or the approximate linear relationship between entropy and prediction error) may not hold strictly in practice, empirical evidence (see Figure˜10) demonstrates a consistent and positive correlation between predictive uncertainty and susceptibility, validating the practical utility of this approach.
@@ C.2 CBM Training Strategies
Here, $\ell_{C_{j}}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ denotes the loss function for the $j$-th concept, and $\ell_{Y}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ denotes the loss for target prediction. Our main experiments are conducted using the Independent strategy, and comparative results for the other strategies are provided in Section˜D.1.
@@ C.3 Implementation Details
For concept prediction, we fine-tune InceptionV3 [47], pre-trained on ImageNet [7], on the CUB [54] and AwA2 [58] datasets. The target predictor $f$ is implemented as a single-layer linear classifier, following the standard setup in Koh et al. [19]. All experiments are conducted with three random seeds, and we report the average performance. Training and evaluation are conducted on NVIDIA GeForce RTX 3090 (24GB), RTX A6000 (48GB), and A100 (80GB) GPUs. The code is available at https://github.com/LOG-postech/CBM-Noise.
@@
All additional CBM variants shown in Section˜D.2 are trained on the CUB dataset for 300 epochs using the Adam optimizer [18] with a fixed learning rate of $10^{-4}$, following Vandenhirtz et al. [52]. For the independently trained autoregressive CBM [14], training is scheduled such that two-thirds of the total epochs are allocated to learning the concept predictor $g$, and the remaining one-third to the target predictor $f$, consistent with Vandenhirtz et al. [52]. We employ ResNet-18 [15] as the backbone for CEM, SCBM, and autoregressive CBM, and ResNet-101 for ECBM. The batch size is set to 256 for CEM, SCBM, and autoregressive CBM models, and 64 for ECBM. When applying SAM to these models, we perform a grid search over $\rho\in\{0.01,0.05,0.1\}$, while keeping all other training hyper-parameters aligned with those used for the Adam baseline.
We conduct a targeted hyperparameter search for each label noise mitigation technique evaluated in Section˜D.5, including label smoothing regularization (LSR) [47], symmetric cross-entropy learning (SL) [56], and SAM. For each method, we tune the corresponding parameters (see Table˜4), such as the smoothing factor $\epsilon$ for LSR, regularization coefficients $\alpha$ and $\beta$ for SL, and the sharpness radius $\rho$ for SAM, under varying levels of noise ($\gamma\in\{0.0,0.1,0.2,0.3,0.4\}$). The best hyperparameters are selected based on task accuracy.
@@ C.4 Label Noise Injection
    Symmetric noise. Each class label is randomly flipped to one of the remaining N1N-1 classes with uniform probability. Specifically, for a dataset with $N$ classes, the probability of mislabeling a sample as any incorrect class is $\frac{1}{N-1}$. This simulates class-independent, uniformly distributed noise. An analogous strategy is applied to concepts, where each concept label is flipped uniformly to a different one.
@@
Symmetric noise captures general annotation uncertainty, asymmetric noise models systematic labeling biases, and grouped noise reflects domain-driven confusions within related concept groups. Together, these diverse noise settings constitute a comprehensive testbed for assessing the robustness of concept-based models. Experimental results under these conditions are presented in Section˜D.4.
@@ C.5 Concept Selection Criteria
This section provides a detailed explanation of the concept selection criteria, extensively analyzed in previous work [43]. We present three criteria (Random, UCP, CCTP) addressed in the main paper, followed by additional criteria (LCP, ECTP, EUDTP) discussed in Section˜D.8.
@@
[…]
@@
[…]
@@ D.1 Effect of Training Strategies under Noise
CBMs can be optimized with three training paradigms: independent, sequential, and joint (see Section˜C.2 for details). We evaluate their robustness by reporting task‑level and concept‑level accuracy across different noise ratios $\gamma$. Figure˜12 presents task accuracy across training strategies, with accompanying radar plots visualizing noise tolerance; larger and more regular polygons reflect greater robustness to noise.
Every training paradigm suffers noticeable degradation as label corruption intensifies. On AwA2, for instance, the joint model starts with the best clean accuracy but still drops from 88.8% to 81.7% when 40% of the labels are noisy. The independent and sequential variants fare even worse and largely collapse on CUB under the same noise level. Table˜5 further reveals declines at the concept level. Although the joint paradigm better preserves task‑level performance, its concept accuracy plunges below 75% once the noise ratio rises above roughly 30% on CUB and exceeds about 10% on AwA2. Because the joint objective simultaneously minimizes task and concept losses, it compromises concept fidelity as label quality declines. These findings corroborate earlier evidence of a performance–interpretability tradeoff [36] and underscore the need for training methods that preserve the robustness of CBM under noisy supervision.
@@ D.2 Robustness of CBM Variants under Noisy Supervision
We also investigate the effects of noise on several CBM variants, including CEM [61], ECBM [59], autoregressive CBM [14], and SCBM [52], trained on the CUB dataset with noise levels $\gamma\in[0.0,0.4]$. Figure˜13 presents task accuracy, while Table˜6 reports concept prediction accuracy under both the standard Adam optimizer (Base) and sharpness-aware minimization (SAM).
@@
D.3 Additional Evidence of Disrupted Representation Alignment
(a) Brewer Blackbird
(b) Importance shift in key concepts
(c) Key concepts under 40% noise
In Section˜3.2, we illustrated how noise disrupts the representation alignment for the class Le Conte Sparrow. Here, we provide additional qualitative results for the class Brewer Blackbird, which demonstrates a comparable degradation pattern.
As shown in Figure˜14, at a moderate noise level (i.e., 20%), the relative ordering of predictive dimensions remains quite well preserved. In contrast, at higher noise levels (i.e., 40%), the saliency of informative dimensions deteriorates significantly, while irrelevant or spurious concepts become disproportionately influential. These findings further substantiate our conclusion that noise induces representational misalignment, and when this misalignment coincides with susceptible concepts, it becomes a key driver of instability in CBMs under noisy supervision.
(a) Task accuracy
(b) Concept accuracy
D.4 Robustness under Diverse Noise Scenarios
We further investigate CBM performance under realistic noise scenarios, specifically asymmetric [13] and grouped noise, using the CUB dataset. Asymmetric noise cyclically mislabels categories (e.g., confusing similar animal species), while grouped noise randomly corrupts labels within similar concept groups (e.g., wing color) as discussed in Section˜C.4.
As shown in Figure˜15, CBMs experience substantial performance degradation under both noise types, broadly consistent with trends observed under symmetric noise. While both asymmetric and grouped noise lead to slightly more pronounced declines in concept accuracy at higher corruption levels (e.g., 40%), the overall performance patterns remain comparable. These findings indicate that CBMs are generally vulnerable to various forms of label noise, reinforcing the need for robustness strategies that extend beyond simple symmetric corruption and account for more realistic noise scenarios.
Furthermore, beyond introducing noise independently, we also considered more practical scenarios where concept and target noise can be correlated. To examine this case, we simulated a correlated noise model and evaluated the relationship between susceptibility and uncertainty. Specifically, we first corrupted the target labels, and concept noise was injected only when the target label was flipped, thereby inducing dependence between concept and target corruption. As illustrated in Figure˜16, the positive relationship between uncertainty and susceptibility remains evident for susceptible concepts. This trend is consistent across different noise levels, suggesting that our analysis and mitigation approach remain effective under more realistic correlated noise conditions.
(a) $\gamma=0.1$
(b) $\gamma=0.2$
(c) $\gamma=0.3$
Figure 16: Correlation between predictive uncertainty and susceptibility on CUB under a realistic conditional noise model. The target label is corrupted first, and concept labels are corrupted only when the target has been perturbed. Pearson correlation coefficients ($\rho$) are reported separately for susceptible and non-susceptible sets, showing a positive correlation in the former.
D.5 Comparative Study of Noise Mitigation Methods
@@
[…]
Table˜7 summarizes task and concept accuracy across noise levels. Across all conditions, SAM consistently outperforms LSR and SL. Under clean supervision, SAM achieves a task accuracy of 79.7%, surpassing LSR (72.8%) and SL (70.6%). Its advantages are more pronounced under noise: at 20% and 30% corruption, SAM outperforms the second-best method by 6.1% and 8.4% in concept accuracy, respectively. These results demonstrate the robustness of SAM in noisy settings. While our evaluation focuses on selected methods, other techniques (e.g., co-teaching [48] or MentorNet [17]) offer promising alternatives through sample selection and teacher-student training. Integrating such techniques with CBMs represents a valuable direction for future work. Overall, these results position SAM as an effective and scalable approach for robust concept-based learning, as further supported by our theoretical insights in Appendix˜A.
D.6 Evaluating Concept Noise in LLM-Generated Annotations